OpenStudy (anonymous):

Which equation best represents the following situation? A 10 pound box of chocolates from Chocolate to Go contains a mixture of their $2.50 per pound Dark Chocolate and their 3.25 per pound cashew cluster. How many pounds of each will be in the box if it cost $16?

OpenStudy (anonymous):

2.50x + 3.25(10 - x) = 16 2.50(10) + 3.25x = 16 2.50x + 3.25(10) = 16 2.50(x) + 3.25(16) = 10x

OpenStudy (anonymous):

If the box costs $16, would that be what the equation costs?

OpenStudy (anonymous):

*equals

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct. so that means that we eliminated the answer to the first three options. Do you get that?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Also, the question says that the box is a 10 pound box of chocolates, but are cashew clusters chocolates?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Is it b?

OpenStudy (anonymous):

EXACTLY. So keeping that in mind and the fact that the chocolates are 2.50 per pound and the clusters are 3.25 per pound, which equation do you think is correct?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

YES!!! Great job!

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

NP! :)